OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −1 A x = 0 B x = 2 C There is no x value as there is no solution. D x can be any value as there is an infinite number of solutions. My answer is B, am I Correct

OpenStudy (anonymous):

rewrite an equation 2x+y=1 y=-2x+1 sub it in the other equation 4x+2(-2x+1)=-1 4x-4x+2=-1 4x-4x=-3 The x's cancel out so there is no other x.

OpenStudy (anonymous):

so i wasn't correct?

OpenStudy (anonymous):

so it would be C right?

OpenStudy (anonymous):

Yes, B and D will not be the answers

OpenStudy (anonymous):

Most likely A or C.

OpenStudy (anonymous):

thanks for your help i say C is the answer :)

OpenStudy (anonymous):

Your Welcome